Find dy/dx if 2y^3t+t^3y=1 and dt/dx=1/cost
1
Expert's answer
2020-05-11T17:55:22-0400

Given "2y^{3t}+t^{3y}=1" and "\frac{dt}{dx}=\frac{1}{cost}".

Let be "F(y;t)=2y^{3t}+t^{3y}-1"

So "\frac{dy}{dt}=-\frac{\frac{dF}{dt}}{\frac{dF}{dy}}" . (1)

"\frac{dF}{dt}=2y^{3t}*3lny+3yt^{3y-1}"

"\frac{dF}{dt}=6y^{3t}lny+3yt^{3y-1}" (2)


"\frac{dF}{dy}=2*3ty^{3t-1}+3t^{3y}lnt"

"\frac{dF}{dy}=6ty^{3t-1}+3t^{3y}lnt" (3)

Substitute (2) and (3) into (1).

"\frac{dy}{dt}=-\frac{6y^{3t}lny+3yt^{3y-1}}{6ty^{3t-1}+3t^{3y}lnt}" (4)


Let's find "\frac {dy}{dx}":

"\frac {dy}{dx}=\frac{dy}{dt}*\frac{dt}{dx}", (5)

where "\frac{dt}{dx}=\frac{1}{cost}". (6)


Substitute (4) and (6) into (5):

"\frac {dy}{dx}=-\frac{6y^{3t}lny+3yt^{3y-1}}{6ty^{3t-1}+3t^{3y}lnt}*\frac{1}{cost}"

"\frac {dy}{dx}=-\frac{6y^{3t}lny+3yt^{3y-1}}{cost(6ty^{3t-1}+3t^{3y}lnt)}" (7)


Answer: "\frac {dy}{dx}=-\frac{6y^{3t}lny+3yt^{3y-1}}{cost(6ty^{3t-1}+3t^{3y}lnt)}".